In the laboratory you dissolve 17.7 g of ammonium acetate in a volumetric flask and add water to a total volume of 125 mL.
1
Expert's answer
2018-04-16T11:07:19-0400
Weight of ammonium acetate = 17.7g
No.mole of ammonium acetate = 17.7/77.082 = 0.23
Total volume = 125 ml
Molarity of solution = 0.23×1000/125 = 1.8 M